# KnobGen: Controlling the Sophistication of Artwork in Sketch-Based Diffusion Models

### v2

@@ Appendix C More Qualitative Result
C.1 Inference Knob Mechanism For Baselines
@@
While the Knob system is designed to balance coarse and fine-grained controls dynamically, the lack of a dedicated coarse-grained module in T2I-Adapter causes the model to lose spatial coherence when we apply our Knob mechanism for it, especially when the knob has low value. This issue became particularly evident when trying to generate images based on prompt only, as the model struggled to infer the missing spatial structure, leading to distorted or incoherent outputs.
@@
C.2 More Qualitative Results
In this section, we present additional qualitative results to demonstrate the effectiveness and versatility of our proposed KnobGen framework further. Figure 10 and  11 showcases the model’s ability to handle a wide range of input sketches, from highly detailed professional-grade drawings to rough, amateur sketches.

### v3

@@ 9 More Qualitative Result
9.1 Inference Knob Mechanism For Baselines
@@
While the Knob system is designed to balance coarse and fine-grained controls dynamically, the lack of a dedicated coarse-grained module in T2I-Adapter causes the model to lose spatial coherence when we apply our Knob mechanism for it, especially when the knob has low value. This issue became particularly evident when trying to generate images based on prompt only, as the model struggled to infer the missing spatial structure, leading to incoherent outputs.
@@
9.2 More Qualitative Results
In this section, we present additional qualitative results to demonstrate the effectiveness and versatility of our proposed KnobGen framework further. Figure 11 showcases the model’s ability to handle a wide range of input sketches, from highly detailed professional-grade drawings to rough, amateur sketches.